# Evaluating Deduplication Techniques for Economic Research Paper Titles with a Focus on Semantic Similarity using NLP and LLMs

### v2

@@ 1 Introduction
The World Bank is exploring the use of causal AI models for analyzing the impact of economic interventions. These models rely heavily on high-quality, accurate data, often derived from academic publications. However, large-scale NLP datasets can be susceptible to the presence of duplicate entries arising from: 1) Title variations: Authors may publish the same research with slightly different titles due to minor wording changes or phrasing variations. 2)Accidental duplicates: Duplicate submissions across academic databases can occur due to human error or technical glitches Christen et al., 2012[Chr12]. 3)Near duplicates: Some research may be published with substantial content overlap but slightly different titles, requiring more sophisticated techniques for identification Zhu et al., 2014[ZG14],Liu et al., 2023[LDL+23], and Elmagarmid et al., 2007[EW07]. Duplicate entries can negatively impact the functionality and accuracy of AI models. This study investigates various deduplication techniques using only titles as a starting point for building a clean and accurate dataset of RCT studies.
@@ 2 Background
Researchers have developed various techniques for deduplication in large NLP datasets, broadly categorized as follows: 1) String-based methods: These techniques rely on string similarity measures like Levenshtein distance Levenshtein, 1966[Lev66] or Jaccard similarity to identify entries with high textual overlap Gomaa et al., 2013[GF13]. 2) Hashing methods: Hashing algorithms map similar textual data points to the same hash value, allowing for efficient duplicate detection. Locality-Sensitive Hashing (LSH) is a popular choice for large datasets Datar et al., 2004[DIIP04] 3) Embedding-based methods: These methods represent text data as numerical vectors using techniques like word embeddings or document embeddings. Similar entries will have vectors close in distance within the embedding space Mikolov et al., 2013[MCCD13], Di Liello et al., 2022[DLGSM22].
@@ 3 Method
    This method calculates the similarity between all possible pairs of titles within the dataset. The computational complexity is given by $\frac{n(n-1)}{2}$, where $n$ is the number of titles Hernández-Caro et al., 2015[HCVM15].
@@
    This method considers titles with similar word counts to be more likely duplicates. We define a threshold ($\delta=5$) for the difference in word count between two titles Sultan et al., 2019[SJG19]. Titles with word count difference less than or equal to $\delta$ are considered for further comparison.
@@
    This method combines source information, word count ranges, and filtering based on the most common word length ($\mu$) in the dataset. Titles from different sources with word counts within a predefined range ($\lambda$) around $\mu$ are selected for comparison Sultan et al., 2019[SJG19]. This condition mathematically as:
@@
    This approach focuses on titles with a length less than or equal to a predefined threshold ($\tau=3$), assuming simpler language and a higher chance of containing duplicates Sultan et al., 2019[SJG19]. Mathematically, this can be expressed as:
@@ 3.1 Distance Measures
    This metric calculates the minimum number of single-character edits (insertions, deletions, substitutions) needed to transform one string (title) into another Levenshtein, 1966[Lev66].
@@
    This method leverages a pre-trained model (SBERT) to generate semantic embeddings for the titles. The distance between the embeddings is used as a measure of similarity Di Liello et al., 2022[DLGSM22]. SBERT models are trained on large datasets of text and sentence pairs, allowing them to capture semantic relationships between words and phrases.

### v3

@@ 1 Introduction
The World Bank is exploring the use of causal AI models for analyzing the impact of economic interventions. These models rely heavily on high-quality, accurate data, often derived from academic publications. However, large-scale NLP datasets can be susceptible to the presence of duplicate entries arising from: 1) Title variations: Authors may publish the same research with slightly different titles due to minor wording changes or phrasing variations. 2)Accidental duplicates: Duplicate submissions across academic databases can occur due to human error or technical glitches Christen et al., 2012[Chr12]. 3)Near duplicates: Some research may be published with substantial content overlap but slightly different titles, requiring more sophisticated techniques for identification Chevallier et al., 2022[CRB+22],Shi et al., 2025[SLL+25], and Dhivyabharathi et al., 2016[VK16]. Duplicate entries can negatively impact the functionality and accuracy of AI models. This study investigates various deduplication techniques using only titles as a starting point for building a clean and accurate dataset of RCT studies.
@@ 2 Background
Researchers have developed various techniques for deduplication in large NLP datasets, broadly categorized as follows: 1) String-based methods: These techniques rely on string similarity measures like Levenshtein distance Miller, 2009[MVM09] or Jaccard similarity to identify entries with high textual overlap Gomaa et al., 2013[WHG13]. 2) Hashing methods: Hashing algorithms map similar textual data points to the same hash value, allowing for efficient duplicate detection. Locality-Sensitive Hashing (LSH) is a popular choice for large datasets Dasgupta et al., 2011[DKS11] 3) Embedding-based methods: These methods represent text data as numerical vectors using techniques like word embeddings or document embeddings. Similar entries will have vectors close in distance within the embedding space Mikolov et al., 2013[MCCD13], Di Liello et al., 2022[LGSM22].
@@ 3 Method
    This method calculates the similarity between all possible pairs of titles within the dataset. The computational complexity is given by $\frac{n(n-1)}{2}$, where $n$ is the number of titles.
@@
    This method considers titles with similar word counts to be more likely duplicates. We define a threshold ($\delta=5$) for the difference in word count between two titles Aronovich et al., 2009[AAB+09]. Titles with word count difference less than or equal to $\delta$ are considered for further comparison.
@@
    This method combines source information, word count ranges, and filtering based on the most common word length ($\mu$) in the dataset. Titles from different sources with word counts within a predefined range ($\lambda$) around $\mu$ are selected for comparison Aronovich et al., 2009[AAB+09]. This condition mathematically as:
@@
    This approach focuses on titles with a length less than or equal to a predefined threshold ($\tau=3$), assuming simpler language and a higher chance of containing duplicates Aronovich et al., 2009[AAB+09]. Mathematically, this can be expressed as:
@@ 3.1 Distance Measures
    This metric calculates the minimum number of single-character edits (insertions, deletions, substitutions) needed to transform one string (title) into another Miller, 2009[MVM09].
@@
    This method leverages a pre-trained model (SBERT) to generate semantic embeddings for the titles. The distance between the embeddings is used as a measure of similarity Di Liello et al., 2022[LGSM22]. SBERT models are trained on large datasets of text and sentence pairs, allowing them to capture semantic relationships between words and phrases.